OpenStudy (waheguru):

How do I know the restrictions on this log function? I need to find its solutions which are... 6 and -2 but there is only one due to restrictions, but which one?

OpenStudy (waheguru):

\[\log_7(x-1)+\log_7(x-5)\]

OpenStudy (turingtest):

what is the domain of \(\log u\) ?

OpenStudy (waheguru):

I dont know

OpenStudy (waheguru):

For tha base function is more than 0

OpenStudy (turingtest):

what does \(\log_by=x\) mean? it means that \(b^x=y\), which means that if \(b>0\) and \(y<0\) there is no solution so \(y\) (the thing you are taking the log of, aka the 'argument') must be greater than zero which of your answers causes one of the arguments to be less than zero?

OpenStudy (waheguru):

6

OpenStudy (waheguru):

I think I get it now...

OpenStudy (turingtest):

let's see\[\log_7(6-1)+\log_7(6-5)=\log_7(5)+\log_7(1)\]none of the arguments are less than zero, so this one is ok

OpenStudy (turingtest):

try the other one

OpenStudy (waheguru):

it doesnt work

OpenStudy (turingtest):

correct\[\log_7(-2-1)+\log_7(-2-5)=\log_7(-3)+\log_7(-3)=undefined\]

OpenStudy (turingtest):

correction:\[\log_7(-2-1)+\log_7(-2-5)=\log_7(-3)+\log_7(-7)=undefined\]

OpenStudy (waheguru):

thanks

OpenStudy (turingtest):

welcome!

OpenStudy (waheguru):

@TuringTest what if both solutions were negative

OpenStudy (waheguru):

or both solutions positive are those senrios possible

OpenStudy (turingtest):

it doesn't matter if the *solutions* are negative, what matters is that they don't make any of the *arguments* (the parts inside the log function) negative

OpenStudy (waheguru):

but are those cases possible

OpenStudy (turingtest):

any number of positive or negative solutions is possible, yes

OpenStudy (turingtest):

it depends on the equation you start with; if it is a 5th order equation you could get 5 different solutions, all of which may be negative or positive or whatever.

OpenStudy (waheguru):

interesting... thanks

OpenStudy (turingtest):

sure thing!